the area of a rectangle is 24 square inches. the lenght of the rectangle is two inches longer tha its width. Find the width of the rectangles
Answer by checkley79(3341)

a=lw
l=w+2
24=(w+2)w
24=w^2+2w
w^2+2w-24=0
(w+6)(w-4)=0
w-4=0
w=4 inches.
l=4+2=6 inches for the length.
Proof:
24=6*4
24=24